# Open-Vocabulary Segmentation with Unpaired Mask-Text Supervision

### v1

@@ Appendix A Framework details
Appendix B Promptable segmentation.
Evaluation details. We perform a prompt segmentation evaluation on a wide range of datasets in various domains. For the point prompt, we adopt a uniform point grid $h\times w$ as input prompts (e.g., 20 $\times$ 20). For the box prompt, we use ground-truth bounding boxes as input prompts. 1-pt IoU denotes the oracle performance of one point by evaluating the intersection-over-union (IoU) of the predicted masks that best match ground truth. 1-box IoU denotes is similar to 1-pt IoU. More evaluation results are reported in Fig. 7, Fig. 8 and Fig. 9.
Dataset details. A description of each dataset is given in Tab. 5. The iShape dataset has 6 subsets, including antenna, branch, fence, hanger, log and wire.
Appendix C Visualisation
We illustrate a wide range of visualisations of promptable segmentation and open-vocabulary segmentation across multiple datasets.
Figure 10: Visualisation of open-vocabulary segmentation between the baseline and ours Uni-OVSeg.
Figure 11: Visualisation of open-vocabulary segmentation between the baseline and ours Uni-OVSeg.
Figure 12: Visualisation of open-vocabulary segmentation between the baseline and ours Uni-OVSeg.
Figure 13: Visualisation of promptable segmentation between SAM-ViT/L and ours Uni-OVSeg.
Figure 14: Visualisation of promptable segmentation between SAM-ViT/L and ours Uni-OVSeg.
Figure 15: Visualisation of promptable segmentation between SAM-ViT/L and ours Uni-OVSeg.
Figure 16: Visualisation of promptable segmentation between SAM-ViT/L and ours Uni-OVSeg.
Figure 17: Visualisation of promptable segmentation between SAM-ViT/L and ours Uni-OVSeg.

### v2

@@ Appendix C Promptable segmentation
Dataset details. A description of each dataset is given in Table 10. The iShape dataset has 6 subsets, including antenna, branch, fence, hanger, log, and wire.
Appendix D Broader impacts
Our approach represents a significant advancement in open-vocabulary segmentation by reducing the reliance on labour-intensive image-mask-text triplet annotations. This innovation has the potential to make cutting-edge vision perception systems more accessible, offering significant benefits across various sectors, such as environmental monitoring and autonomous vehicles. Developing a more efficient and accurate vision perception system contributes to the broader field of AI, potentially leading to more innovative applications and research in machine learning, computer vision, and related areas. As with any AI model, the risk of bias in the data used for training is a crucial concern. Efforts must be made to ensure the datasets are diverse and representative to avoid perpetuating or amplifying biases.
Appendix E Visualisation
We illustrate a wide range of visualisations of prompt segmentation and open-vocabulary segmentation across multiple datasets.
Figure 11: Visualisation of open-vocabulary segmentation between the baseline and Unpair-Seg.
Figure 12: Visualisation of open-vocabulary segmentation between the baseline and Unpair-Seg.
Figure 13: Visualisation of open-vocabulary segmentation between the baseline and Unpair-Seg.
Figure 14: Visualisation of promptable segmentation between SAM-ViT/L and Unpair-Seg.
Figure 15: Visualisation of promptable segmentation between SAM-ViT/L and Unpair-Seg.
Figure 16: Visualisation of promptable segmentation between SAM-ViT/L and Unpair-Seg.
Figure 17: Visualisation of promptable segmentation between SAM-ViT/L and Unpair-Seg.
Figure 18: Visualisation of promptable segmentation between SAM-ViT/L and Unpair-Seg.